OpenStudy (anonymous):

Is this right? Solve this: x-8/5=x/9 my answer: 14x-72/45

sam (.sam.):

x-(8)/(5)=(x)/(9) 9x-(72)/(5)=x 8x-(72)/(5)=0 8x=72/5 x=9/5

OpenStudy (anonymous):

its a rational expression are you sure?

sam (.sam.):

it depend whether \[\frac{x-8}{5}=\frac{x}{9}~~~~OR~~~~x-\frac{8}{5}=\frac{x}{9}\]

OpenStudy (anonymous):

It's the first one

sam (.sam.):

(x-8)/(5)=(x)/(9) x-8=5x/9 9x-72=5x 4x-72=0 4x=72 x=18

OpenStudy (anonymous):

How the heck did you get that? :( I tried but I couldnt get a straight answer

sam (.sam.):

 Cross multiply